A food safety guideline is that the mercury in fish should be below 1 part per million​ (ppm). Listed below are the amounts of mercury​ (ppm) found in tuna sushi sampled at different stores in a major city. Construct a 98​% confidence interval estimate of the mean amount of mercury in the population. Does it appear that there is too much mercury in tuna​ sushi? 0.56  0.82  0.10  0.96  1.33  0.53  0.89 What is the confidence interval estimate of the population mean u?​
Answer #1

solution: n=7 0.56 0,82 0.10 0.96 1:33 0,53 0. 89 5.19 0.3136 0,6724 0.01 0.9216 1.7689 | 0.2809 0.7921 4.76 Sum= mean i txi,is At 98% confidence interval t C = 98% = 0.98 = 1-C = -0.98 = 0.02 Ź 29,02 - 0.01 d. f. = n-1 = 7-1 - 6 42, d.f. = 8.01,6 =